The marketing research department for a company that manufactures and sells laptop computers

established the following price-demand and cost functions, respectively:

p (x) = 1,000 - 25x

C(x) = 2,150 + 250x

where p() is the wholesale price in United States dollars at which y thousand laptops can be sold and

C(x) is in thousands of dollars. The domain of both functions is, 1 < x < 10.


formulate a revenue and a profit function


1
Expert's answer
2022-02-27T02:18:35-0500

The revenue function is given as : "R(x)=P(x)\u00d7x"

"R(x)=(1000-25x)x"

"R(x)=1000x-25x^2"


The profit function, "\pi(x)", is given as: "\pi(x)=R(x)-C(x)"

"\pi(x)=1000x-25x^2-(2150+250x)"

"\pi(x)=1000x-25x^2-2150-250x"

"\pi(x)=750x-25x^2-2150"